OpenStudy (anonymous):

A rectangle has an area of x3 + 5x2 + 5x – 2 square meters and a width of x + 2 meters. Find its length.

OpenStudy (mathstudent55):

Do you know the formula for the area of a rectangle?

OpenStudy (anonymous):

w*h?

OpenStudy (mathstudent55):

Right, but it will be easier if we use the terms the problem uses. Instead of dealing with a width and a height, we will deal with a length and a width. The area is: A = L * W

OpenStudy (anonymous):

These are my choices, x2 + 5x – 4 meters x2 + 7x – 9 meters x2 +3x – 1 meters x2 + 3x + 1 meters

OpenStudy (mathstudent55):

The problem gives us the area and the width, and it asks for the length. Let's solve the formula for the length, what the problems is asking for. \(A = L \times W\) We divide both sides by W to isolate L: \(\dfrac{A}{W} = L\) or \(L = \dfrac{A}{W} \)

OpenStudy (mathstudent55):

Now we need to do the division: \(A = \dfrac{x^3 + 5x^2 + 5x – 2}{x + 2} \) Now you can do long division of polynomials to find the answer.

OpenStudy (mathstudent55):